OpenStudy (anonymous):

a. Use a metric ruler to measure the distances between Robinson and Neale on the map. b. Using the scale of the map, find the approximate actual distance by air (not by roads), between Robinson and Neale. c. Approximately how many square miles are shown on this map?

OpenStudy (anonymous):

http://www.intervisualtechnology.us/data/1/12383/texasmap.jpg is the link to the map

OpenStudy (misssunshinexxoxo):

Ok please pull out your ruler and mark it down then post it; I'll go over it

OpenStudy (misssunshinexxoxo):

2.4 centimeters is what you messaged me; what's the mileage

OpenStudy (misssunshinexxoxo):

Looks like 4 miles to me

OpenStudy (anonymous):

Thats the part i'm having trouble with. The scale says 0-5 miles, but idk how to incorporate that with the 2.4 centimeters.

OpenStudy (misssunshinexxoxo):

Is this an essay question?

OpenStudy (anonymous):

no

OpenStudy (misssunshinexxoxo):

I think you separate it

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

So you measured the distance between Robinson and Neale and got 2.4 cm?

OpenStudy (anonymous):

yea

jimthompson5910 (jim_thompson5910):

use a ruler to measure how long the "0-5" bar is

OpenStudy (anonymous):

2.1 cm

jimthompson5910 (jim_thompson5910):

x = actual distance from Robinson to Neale solve the proportion for x \[\Large \frac{\text{Actual Distance}}{\text{Distance on paper}} = \frac{\text{Actual Distance}}{\text{Distance on paper}}\] \[\Large \frac{x}{2.4} = \frac{5}{2.1}\] \[\Large x = ???\]

OpenStudy (anonymous):

5.7

OpenStudy (anonymous):

but it says to find the distance by air and not by roads, what about that part?

jimthompson5910 (jim_thompson5910):

that's the actual air distance

OpenStudy (anonymous):

and it would be 5.7 for part C also?

jimthompson5910 (jim_thompson5910):

measure the width and height of the entire map with the ruler

jimthompson5910 (jim_thompson5910):

what do you get?

OpenStudy (anonymous):

6.8 cm length and width

OpenStudy (anonymous):

so do i just multiply that now?

jimthompson5910 (jim_thompson5910):

well you first have to convert those distances on paper to actual distances

jimthompson5910 (jim_thompson5910):

by using a proportion like shown above

OpenStudy (anonymous):

so then 16.19

OpenStudy (anonymous):

i mean 16.19 is the length and width, and multiplying it i get 262.12

jimthompson5910 (jim_thompson5910):

so you just found that 6.8 cm on paper corresponds to 16.19 miles in actual distance

jimthompson5910 (jim_thompson5910):

and yeah the approximate actual area is 262.1161 square miles

OpenStudy (anonymous):

yea, i set it up like 5/2.1 = x/6.8 is that right?

jimthompson5910 (jim_thompson5910):

yes it is

OpenStudy (anonymous):

cool, thank you!

jimthompson5910 (jim_thompson5910):

no problem